An echo is heard from a building 0.721 s after you shout "hello." How many feet away is the building if the air temperature is 21.5°C?
1
Expert's answer
2019-04-11T09:35:20-0400

Let's first find the speed of sound at temperature 21.5°C from the formula:


"c = \sqrt{\dfrac{\gamma R T}{M}},"

here, "c" is the speed of sound at temperature 21.5°C, "\gamma = 1.402" is the adiabatic index of air, "R = 8.3145 \dfrac{J}{mol \cdot K}" is the molar gas constant, "T" is the temperature in Kelvin and "M = 0.0289645 \dfrac{kg}{mol}" is the molar mass of air.

Then, we get:


"c = \sqrt{\dfrac{1.402 \cdot 8.3145 \dfrac{J}{mol \cdot K} \cdot (273.15 + 21.5^{\circ}C)}{0.0289645 \dfrac{kg}{mol}}} = 344.36 \dfrac{m}{s}."

Then, we can find the distance from the building to the person from the formula:


"c = \dfrac{Total Distance}{Time} = \dfrac{2s}{t},"

here, "s" is the distance traveled by the sound to the building (since the sound reflects back from the building to the person, the total distance traveled by the sound is doubled), "t" is the time.

Therefore, from this formula we can find the distance from the building to the person:


"s = \dfrac{1}{2} ct = \dfrac{1}{2} \cdot 344.36 \dfrac{m}{s} \cdot 0.721 s = 124.14 m \cdot \dfrac{1 ft}{0.3048 m} = 407.3 ft."

Answer:

"s = 407.3 ft."